2001-02 Official Log

From OCE Space Simulation
Jump to: navigation, search

The Mission

Mission to 
Mars (Mars 2002)

The Log

T-00:20:00 Countdown begins.
T-00:00:00 Liftoff.
T+00:04:35 SRB sep; negative return.
T+00:09:07 External tank separation.
T+00:10:07 Habitat power remotely initialised.
T+00:12:50 Three astronauts have lost consciousness; possible causes include G-forces and heat exhaustion.
T+00:15:14 Astronauts told to expect vibrations on the port side of the shuttle.
T+00:16:18 Astronauts report no turbulence; MC confirms that the danger of turbulence is over.
T+00:17:20 Leigh has regained consciousness; no change in the other two.
T+00:18:28 Astronauts are in the corridor.
T+00:18:57 Kelli collapses; there are now three unconscious astronauts.
T+00:19:30 Kevin speaks and is therefore considered conscious; there are now two unconscious astronauts.
T+00:24:13 All astronauts have regained consciousness except for Kelli.
T+00:26:10 Astronauts are informed that they should all be awake and preferably standing up.
T+00:29:26 Astronauts are told that they should all stand on the Z-axis.
T+00:31:28 Astronauts report hearing a noise; they are informed it was due to space debris.
T+00:35:18 Rendezvous with habitat completed.
T+00:35:58 Umbilical cord pressurised.
T+00:36:59 Astronauts open umbilical door.
T+00:40:25 Astronauts enter the habitat.
T+00:41:43 Astronauts begin to power up the hab.
T+00:43:03 Astronauts report that the bathroom is the only module with air circulation.
T+00:44:25 The auxiliary camera is online.
T+00:45:16 Main camera system is online.
T+00:45:44 Radiators are stowed.
T+00:48:29 Astronauts told to test their stereo equipment.
T+00:50:23 Camera systems being worked on; astros told to ensure that their TVs are plugged in.
T+00:51:19 MC is not getting signals from either Habitat TV; the problem is put on hold for resolution at a later time.
T+00:53:45 Shuttle has re-entered the atmosphere.
T+00:55:42 Habitat has reached lunar orbit.
T+00:56:20 AYSE drive begins to dock with the habitat.
T+01:00:19 AYSE drive docking complete.
T+01:05:33 Habitat has initialised GNC route.
T+01:11:30 Stereo test has begun; work on cameras continues.
T+01:13:07 Shuttle is not falling fast enough.
T+01:19:20 Shuttle is in the pipe 5x5; pilot believes he can land safely.
T+01:22:58 Shuttle has touched down.
T+01:32:30 Official mission clock is back up and running.
T+01:36L40, Astronauts request an emergency supply probe. More items added for remainder of day.
T+01:40:36 Leigh is now the habitat CapCom.
T+01:49:56 Habitat ventilation system is down.
T+02:00:00 Mission clock appears to be no longer working.
T+02:00:50 Roman's watch designated new standard for measuring mission time. Asume mission began at 15:50.
T+02:03:34 Alex waves a light in front of one of the Hab cameras.
T+02:09:01 Roman decides that previous times were all twenty minutes ahead. Log times adjusted accordingly.
T+02:10:20 Mission time counter is officially working again. Resume taking time from this counter.
T+03:40:20. Emergency Probe Launched.
T+15:59:44 Logs re-opened, for beginning of Day 2.
T+16:01:18 Cap Com is not working. Ross Green and Roman attempt to fix the problem.
T+16:15:20 Astronauts iniate emergency status. Hab powered down, while ventillation errors are repaired and one outlet is repaired.
T+16:51:29 Astronauts are told to open a set of sealed orders, and prepare for an EVA in approx. one hour.
T+17:36:50 Astronauts told to prepare for EVA at 10:30 (mission time 18:30).
T+17:40:00 MC experiences problems with AuxCom; requests instructions for re-starting programme from Hab.
T+17:43:49 Hab suggests switching AuxCom and Logs computers.
T+17:45:00 login for hyperbola requested
T+17:46:58 AuxCom back in service
T+17:47:47 request for information on conneting the transformer
T+17:52:46 astraunots reminded of the EVA in 15 minutes
T+17:53:34 Alex and Leigh are the astronauts who are to go out on the surface
T+17:54:07 Astronauts going about their business
T+17:55:08 Astronauts informend of the time they have before they come outside of the habitat
T+17:56:00 Fuses for Robbie requested
T+17:56:06 Request from Kellie for Roman to get into his flight suit
T+17:56:09 Roman obeys orders
T+17:56:11 Roman reminds astronauts have 7 minutes until Eva
T+17:56:12 Astros confirm ETA 7 minutes
T+18:15:56 Astros request new ETA
T+18:16:54 new ETA is supplied
T+18:17:30 Astronauts are asked to confirm they will be proceeding with EVA
T+18:18:22 Alex told to give Leigh the headset because Alex is unintellegible
T+18:19:25 Leigh has the EVA capcom
T+18:20:14 Roman asks for a picture taking device,reminds astros to take one
T+18;21:05 Roman checks the time of day for the astros
T+18:22:03 they are told that it is night time - light -minimum
T+18:24:15 astros 5 minutes behind EVA schedule
T+18:26:06 Astronauts supplying EVA with Duct Tape for repair.
T+18:27:44 Leigh testing headset
T+18:28:48 Requesting confirmation for depressurization of airlock
T+18:32:55 Airlock depressurized
T+18:34:10 Requwst from mission control to view picture from the habitat
T+18:35:06 Request for Sit Rep
T+18:35:36 Mission control asks for confirmation that intel cam is connected where it is supposed to be
T+18:37:28 Answer inaudible - confirmation
T+18:39:50 Astros report on surface
T+18:40:18 Astros told to look for anyhting savageable of the electornics
T+18:41:23 Astronauts report of keybord and speakers found, keybord the wrong type
T+18:42:40, Astronauts report on an 'unstable' cave behind the Habitat. Rocks always shifting inside cave.
   Sometimes blocked,sometimes unblocked.
T+18:43:45 Astronaut reentering airlock
T+18:52:40 Airlock pressurized
T+18:53:20 Astronauts back in habitat
T+18:55:52 Astronauts report not to have experienced any metereological problems
T+18:56:00 Report from astronauts soon to arrive
T+18:57:54 Astros asked if they have found a CD
T+18:59:54 they have found one
T+19:01:55 Alex requests to speak with mission commander
T+19:02:34 Ross is occupied and unable to talk to him
T+19:03:02 Report being sent through AuxCom
T+19:04:24 Mission control says probe will be landing in 16 minutes
T+19:06:20 Roman passes headsets to Wajid
T+19:08:46 Kevin on CapCom
T+19:14:55 Yi Yuan's pre-exercise stats: 144sys 86di p46
T+19:27:22 MC confirms for the astronauts that CMES is down, as is the network on <Pi>ro
T+19:29:50 Kevin reports static on his headset
T+19:51:30 White team goes on shift
T+20:02:00 Bryan takes over logs
T+20:03:09 Overhead is lost
T+20:03:20 Overhead is back
T+20:12:26 Head sets are tested
T+20:13:50 Astronauts are putting on space suits
T+20:16:40 Astronauts are putting on space suits in interlock
T+20:17:20 Alex goes into to a mumbling fit
T+20:19:33 Depressurizing airlock
T+20:19:50 Alex quickly escapes the airlock
T+20:21:35 Bo takes over for Bryan
T+20:25:00 Depressurizing now
T+20:30:42 Airlock is depressurized
T+20:34:40 Repressurizing airlock
T+20:40:25 Repressurizing is complete
T+20:42:44 Helmet broke, 3 left
T+20:47:42 Bryan takes over logs
T+20:58:00 Problem with rear headsets found
T+20:59:01 Bo takes over for Bryan....again
T+21:09:31 Leigh's biomed after cycling: sys205 dias172 pulse68
T+21:10:18 Power in longhouse lost and cams are off
T+21:11:22 Projector video lost
T+21:11:40 Projector is back
T+21:11:46 Power is back in longhouse
T+21:13:17 Source of problem found. A plug seems to be missing
T+21:23:16 Graham is now on Capcom
T+21:26:04 Bryan takes over Auxcom
T+21:30:00 Kelli is on CDO
T+21:33:30 Graham is now on logs
T+21:35:11 Yi Yuan and John are suiting up for an EVA
T+21:41:30 Yi Yuan and John have entered the airlock
T+21:48:53 Sealing airlock
T+21:49:24 Depressurizing airlock
T+21:55:02 Airlock depressurized
T+21:55:24 Astros have left the airlock
T+22:03:14 Astros are returning from EVA because of reports of a dust storm
T+22:04:24 Repressurizing Airlock
T+22:09:13 Airlock repressurized. Astronauts are back in teh airlock.
T+22:11:39 Leigh is now on capcom
T+22:16:34 Dust Storm has hit the hab. All Power in the interlock is out
T+22:18:18 Auxilary Camera system is running
T+22:19:04 Astronauts are attempting to repair the problem
T+22:19:53 Removing EVA suits from airlock
T+22:20:57 Astronauts are moving into longhouse
T+22:23:22 Preparing to Check breaker in power room using EVA suits
T+22:26:23 Astronaut John is checking the power room
T+22:35:50 Astronauts have power restored to entire Habitat
T+22:58:28 Kevin's Rest Biomed Stats: 119, 72, 79.
T+23:24:19 Kevin's After Biking Bios: 128/63 p80.
T+23:43:03 MC attempts to find a solution to our chronic headset shortage.
T+23:48:17 Kelli's post-biking stats: pulse 70 blood pressure 136/61
T+23:51:03 Astronauts begin brainbalancing.
T+23:58:29 Dan is now on logs.
T+23:59:03 Capcom speakers are down for a time now and will continue to be so.
T+24:00:00 One day has passed since the mission commenced.
T+24:00:49 Brainbalancing is continuing.
T+24:04:30 Another headseat is now available in Mission Control <-- Dave fixed one.
T+24:53:44 Ms. McLean has arrived.
T+24:56:15 Request by flight director for cameras to be turned on.
T+24:58:10 Attempting to upgrade Hyperbola to add SSH.
T+25:00:16 Mission Control Capcom speakers are back online.
T+25:20:05 Cameras are online.
T+25:48:08 Dan has taken over Capcom.
T+25:53:58 Alex has lost communication.
T+26:01:01 Astronauts have finally noticed the camera test set up by Ross. They failed
  the test due to their slow reaction time  (over 25 min).
T+26:02:02. Astronauts inform MC they noticed immediately, but did not report because they didn't want to antagonize MC.
T+26:12:12 Adrian took Capcom.
T+27:04:09 Habitat has lost all power; there appears to be smoke in the habitat; possible
    hull breaches due to a meteor shower. All  astronauts safe in the longhouse. Emergency Declared.
T+27:05:13 There is liquid on the electronics.
T+27:06:30 Astronaut Alex goes out on an EVA to determine and correct the problem.
T+27:08:28 There is no blast crater around the habitat; Alex looks for damage to the power source.
T+27:09:44 Alex's headset is losing power. However, there is no visible damage to the habitat.
T+27:10:49 Roman launches in case he is needed to rescue the astronauts.
T+27:12:03 Airlock is pressurised for Alex to re-enter the habitat.
T+27:13:22 Alex's headset is dead.
T+27:14:46 Alex enters the interlock.
T+27:17:10 Alex reports the interlock is full of a powder-like airborne substance and there 
   is a blood-like substance in the power  box. 
T+27:18:49 Roman is in Earth orbit, ready for possible extraction.
T+27:21:39 Alex prepares his equipment for a second IVA to the interlock.
T+27:25:44 Alex and Kevin enter the interlock.
T+27:27:30 Alex returns to the longhouse and switches headsets.
T+27:30:55 The liquid in the power box is believed to originate from within the habitat; it may be juice.
T+27:34:34 Alex reports no sparking in the power box.
T+27:35:05 The liquid is confirmed to be juice.
T+27:38:08 Astronauts now believe the liquid is paint thinner.
T+27:52:43 Astronauts try to mop up the substance and turn on the power.
T+27:58:40 Alex's transmissions are garbled and cannot be understood.
T+27:59:38 Substance in the power box is no longer liquid; it has been partially absorbed by the drywall.
T+28:04:14 Power is restored; permission is given for astronauts to enter the interlock.
T+28:07:53 Tech Director suggests that the liquid may have been lubricating fluid from the gears of the satellite dish.
T+28:08:54 Robbie is declared inoperative: damaged during meteor strike.
T+28:10:03 Longhouse cameras and TV are inoperative.
T+28:11:35 Astronauts repair cameras and TVs.
T+28:16:57 Astronauts are attempting to upgrade Andromeda; auxcom is inoperative. Upgrade interupted 
   by meteor strike. Auxcom destroyed.
T+40:20:41 logs back online, day 3
T+40:21:13 Bo is on logs
T+40:24:49 Airlock monitor back online
T+40:37:08 Kelli's biomed: bloodpressure105/69 pulse62
T+40:47:35 Jing Ping is back on Capcom
T+40:49:55 Leigh's precycling biomed: bldpres105/56 pulse78
T+40:58:46 John is still asleep
T+41:00:57 John is conscious .
T+41:03:12 Nik is in
T+41:08:44 Projector video lost
T+41:12:56 The plant isn't dead yet
T+41:13:31 Kevin & John on biomed
T+41:14:28 Projector is back online
T+41:16:08 Projector is off
T+41:20:23 Graham is on logs
T+41:21:06 Leigh's biomeds: bloodpressure99/62 pulse 79
T+41:22:30 John's biomeds: bloodpressure113/70 pulse 87
T+41:23:23 Kevin's Biomeds: bloodpressure120/75 pulse 96
T+41:24:13 Alex's Biomeds: bloodpressure116/73 pulse 82
T+41:31:36 Yi Yuan's Biomeds:bloodpressure138/83 pulse 71
T+41:34:06 Astronauts Kelli and Yi Yuan are suiting up for an EVA
T+41:47:38 sealing airlock
T+41:49:12 depressurizing airlock
T+41:52:23 airlock depressurized. Astronauts are leaving habitant.
T+41:54:28 Problem with Kelli's headset.
T+41:57:45 A dust storm begins on Mars. Astro's return to airlock.
T+41:58:58 Repressurizing airlock
T+42:02:38 Airlock repressurized.Astronauts are back in the habitat.
T+42:03:11 Internet is back online: website updated.
T+42:06:45 Cameras are down; large amounts of static on CapCom. Emergency Declared.
T+42:09:14 Auxcam shows that there is no power in the interlock.
T+42:10:25 Astronauts report that the power has been cut throughout the habitat.
T+42:12:03 Habitat does not appear to be receiving MC's transmissions.
T+42:12:33 Roman is on station (in Earth orbit).
T+42:12:52 Habitat launches a relay satellite.
T+42:16:58 CapCom is back up; astronauts attempt to re-establish visual contact.
T+42:17:51 Cameras are restored; power is evidently back up.
T+42:22:26 Shift change: Jingping on logs
T+42:24:11 Alex's precycling biomeds: bloodpressure: 144/81 pulse: 87
T+42:48:49 Alex's postcycling biomeds: bloodpressure: 132/64 pulse: 135
T+43:02:08 Shift changes: Jennifer on Logs
T+43:18:49 Yi Yuan's Biomeds - before biking: bp 151/125 pulse 74
T+44:33:48 Analysis of EVA samples: 
   1. two identical rocks with many brittle layers
   2. medium sized grey rock, semi lustrous, very hard and solid
   3. grey sedimentary rock, about same size as #2, has minimal lustre, smooth texture
   4. small red rock, very hard, semi-lustrous, transluscent white substance on it, possible biohazard, possibly radioactive
   5. dust from mars, rough grain, somewhat red with large black particles, somewhat lustrous
T+? Kevin's Biomeds: bp 128/86 pulse 89
T+45:12:35 Since mission time counter is not working, Roman's watch becomes standard.
T+45:15:08 Kevin's biomeds: bp 133/80 pulse 77
T:46:05:00. Alex and Kevin go on reapir and inspection EVA
T+46:21:10. Astronauts return from EVA
T+46:53:39 Kelli's precycling biomeds: bp 128/61 pulse 77
T+46:17:37 Kelli's postcycling biomeds: bp 115/70 pulse 70
T+46:47:33 John's postcycling biomeds: bp 134/76 pulse 92
T+48:48:40 Dust storms are headed towards the habitat.
T+48:51:41 Cameras in interlock are on the fritz; astros turn off electronics. Emergency Declared.
T+48:52:16 Power fluctuations in interlock.
T+48:52:36 Lots of small impacts and large impacts at 10-15s intervals.
T+48:54:47 Longhouse does not appear to be affected by dust storms.
T+48:57:33 Interlock readings deteriorating; loud noises heard.
T+48:59:00 Astronauts ordered to evacuate the interlock.
T+49:00:31 Astronauts re-inforce bulkhead. Bulkhead on North East side of interlock most affected.
T+49:01:00 Astronauts rout power from auxiliary circuit to the auxiliary fan. Additional air now being pumped into habitat.
T+49:01:40 Astronauts report the bulkhead is holding.
T+49:03:42 Wajid's headset does not work; he switches batteries.
T+49:04:16 Astronauts report the bulkhead is secure.
T+49:07:33 Astronauts evacuate the interlock, into bathroom. Leave bulkhead between rooms open to obvserse re-inforced bulkhead.
T+49:09:02 Alex reports that there is no hull breach.
T+49:10:30 Alex reports bulkhead has decompressed explosively. Direct meteor inpact. Astronauts
    close bulkhead from bathroom to   interlock, and retreat to longhouse. 
T+49:12:41 Astronauts hear more collisions, this time in the bathroom area.
T+49:16:38 Interlock sensors are somewhat jumbled; readings show extremely low pressure and oxygen.
T+49:20:43 Astronauts report that the dust storm seems to have quietened.
T+49:25:39 Astronauts report no unusual noises.
T+49:29:08 Astronauts turn on more lights.
T+49:32:22 Authorization granted for Kelli to perform an IVA to the interlock.
T+49:34:20 MC plans to vent air from the shower's air tank as Kelli enters the bathroom; 
  this should prevent explosive  decompression of the longhouse should the bathroom be depressurized.
T+49:36:20 Sensor readings seem to indicate that the bathroom is pressurized; however, air will 
  still be vented as a pre-emptive measure.
T+49:39:17 Astronaut Yi Yuan reports hearing a thud outside the habitat.
T+49:40:07 Static on the headsets.
T+49:42:00 Astronauts are down to one roll of duct tape; concern about Kelli's ability to 
  fix hull breaches with only one roll of  duct tape and no other astros on the IVA.
T+49:53:37 Airlock depressurized so that Kelli can retrieve the EVA suits from it.
T+49:55:07 Kelli enters the bathroom and inspects it quickly for damage.
T+49:56:04 Kelli attempts to clear away debris that had fallen into the interlock.
T+49:57:12 Kelli retrieves the other EVA suit from the airlock.
T+49:58:06 Repressurizing the bathroom.
T+50:03:37 Bathroom fully repressurised; Kelli re-enters the longhouse.
T+50:21:58 Bathroom depressurisation begun for a second IVA (Kelli and John).
T+50:24:04 Bathroom depressurisation complete; astronauts enter interlock.
T+50:29:42 Repairs are halfway completed.
T+50:36:13 Repressurisation of interlock commences.
T+50:40:27 Power and video restored in the interlock; interlock is 80% pressurised.
T+50:42:25 Interlock is fully pressurised.
T+50:43:08 Astronauts re-enter the interlock and desuit Kelli and John.
T+50:48:55 Astronaut Kevin is not feeling well; Leigh believes he is just tired.
T+51:00:55 Confirmation that Hyperbola is running.
T+51:06:16 Bonk shuts down when there was an attempt to run Auxcom
T+51:13:53 Astronauts allowed to have dinner.
T+51:17:53 Bonk back up.
T+51:28:55 Alex: "I'm going to see if I can install Satan on Hyperbola."
T+51:29:40 Alex continues his search for Satan.
T+51:45:30 Power surge causes Stonehenge to temporarily go down.
T+55:38:00 Astronauts report slow leak in Interlock ceiling. Evacuate the Longhouse. Prep for EVA. Emergency Delcared.
T+56:05:00 John repairs leak, Emergency over.
T+56:38:00 Astronauts leave message on AuxCom indicating that they again have a hull breach 
  in the Interlock. This breach is larger  than the first. Astronauts lie down and sleep to conserve oxygen.
T+62:59:30 Logs reactivated for final day.
T+63:00:00. Astronauts retreat to Longhouse. Interlock at 73% oxygen. Astronauts await MC staff to preform EVA.
T+64:45:00 Astronauts decide to preceed with EVA without MC staff.
T+64:55:24 Astronauts Alex and John depart on an IVA for repairs.
T+65:03:47 IVA astronauts return to Longhouse.
T+65:10:11 Interlock repressurized; astronauts re-enter.
T+65:24:30 John pre-cycling: bp 130/74, pulse 78
T+65:35:05 John post-cycling: bp 126/70, pulse 86
T+66:21:10 Yi Yuan pre-cycling: bp 163/78, pulse 95
T+67:04:29 Dan is now on logs
T+67:05:23 Shutting down in order to commence deorbiting procedures
T+67:17:33 Astronauts are entering the shuttle
T+67:18:12 Astronauts are ready.
T+67:19:01 Shuttle Galileo requesting fire-for-effect.
T+67:25:15 Shuttle Galileo is setting to manual.
T+67:26:03 Shuttle Galileo de-orbit burn completed.
T+67:35:05 Shuttle Galileo is in entry phase.
T+67:3------ Mission timer is down
T+67:-------- Preparing for landing procedures.
T+67:-------- Shuttle Galileo has incorrectly prepared the gears.
T+67:-------- Astronauts report environmental systems problem. MC states that pilot is unaware of this and that mission will go as  planned.
T+67:-------- Galileo reports KU antenna is probably fried.
T+67:-------- Shuttle is at 300 knots; too fast
T+67:-------- Galileo is on final approach.
T+67:-------- Undercarriages deployed.
T+67:-------- Galileo passed point-of-no-return.
T+67:-------- Successful touchdown!
T+67:-------- Galileo has reached full stop.
T+67:-------- Astronauts reached home.